In a graduate teacher college, a survey was conducted to determine the proportion of students who want to major in mathematics. If 368 out of 850 students said yes, with 95% confidence, what interpretation can we make regarding the probability that all students in the teacher graduate college want to major in mathematics?


1
Expert's answer
2021-06-02T12:51:07-0400

Here,

X = 368

n = 850


"p'=\dfrac{X}{n}=\dfrac{368}{850}=0.433"


95% confidence interval

"So \ \alpha=1-0.95-0.05"

"\implies Z_{\alpha/2}=1.96"


Now, "q'=1-p'=1-0.43=0.57"


and "\sqrt{\dfrac{p'q'}{n}}=\sqrt{\dfrac{0.43\times 0.57}{850}}=0.0169"


for the lower limit:

"p'-Z_{\alpha/2}\sqrt{\dfrac{p'q'}{n}}=0.433-(1.96)(0.0169)=0.396\ or\ 39.6\%"



for the upper limit:

"p'+Z_{\alpha/2}\sqrt{\dfrac{p'q'}{n}}=0.433+(1.96)(0.0169)=0.464\ or\ 46.4\%"


Thus, with 95% confidence, we can state that the interval from 39.6% to 46.4% contains the true percentage of all graduate students who want to major in Mathematics.